There are 6 performers who will present their comedy acts this weekend. One of the performers insists on being the last stand-up comic of the evening. If this performer's request is granted, how many different ways are there to schedule the appearances?
I tried to use the permutations formula where n=6(total performers) and r=5(other performers to arrange)
I came up with 720 ways.
Am I using the right formula? And steps?
Help please.

Answer by richwmiller(17219)

Since the last place has been assigned the first place has only 5 choices
5*4*3*2*1=20*6=120
not
6*5*4*3*2*1=720